Area of a rectangular field is 120 sq.m and perimeter is 46 m. Findout the dimensions of the field.
Answer by Alan3354(69443)

Area of a rectangular field is 120 sq.m and perimeter is 46 m. Findout the dimensions of the field.
----------------
L*W = 120
L + W = 23 --> W = 23 - L
--------------
L*(23 - L) = 120
23L -L^2 = 120
L^2 - 23L + 120 = 0
(L - 8)*(L - 15) = 0
-------------
--> 8 by 15